OpenStudy (anonymous):

30 x 30(x) = 25 x 55(x) solve for x

OpenStudy (akashdeepdeb):

We can and NEVER SHOULD cancel variables!! So 30*30 * x/25*55 = x 36x = 55x or 55x - 36x = 0 x(19) = 0!! :D So x=0! :)

OpenStudy (anonymous):

thank you but I am sorry, the two equations were addition, my bad. It was supposed to look like this, I am so sorry 30 + 30(x) = 25 + 55(x) solve for x

OpenStudy (akashdeepdeb):

 Haha...Does this help? :)